### v2

# Positive weak solutions of a double-phase variable exponent problem with a fractional-Hardy-type singular potential and superlinear nonlinearity

Abstract
In the present paper, we study a double-phase variable exponent problem which is set up within a variational framework including a singular potential of fractional-Hardy-type. We employ the Mountain-Pass theorem and the strong minimum principle to obtain the existence of at least one nontrivial positive weak solution.
keywords:
Positive weak solution; double phase operator; variable exponents; fractional-Hardy-type singularity; the Mountain-Pass theorem; the Ambrosetti–Rabinowitz condition; strong minimum principle.
articletype: Research Paper
1 Introduction
In this article, we study the following singular double-phase variable exponent problem
$\begin{cases}\begin{array}[]{rlll}-\mathcal{D}(x,\mu,\nabla u)+\mathcal{S}(x,\mu,u)&=\lambda f(x,u)\text{ in }\Omega,\\ u&=0\text{ on }\partial\Omega,\end{array}\end{cases}$ ($\mathcal{P}_{\lambda}$)
where
$\mathcal{D}(x,\mu,\nabla u):=\mathrm{div}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u),$
$\mathcal{S}(x,\mu,u):=\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}.$
We assume that $\Omega$ is a bounded domain in $\mathbb{R}^{N}$ $(N\geq 3)$ with Lipschitz boundary $\partial\Omega$; $f$ is a Caratheodry function; $1<p,q\in C(\overline{\Omega})$; $0\leq\mu(\cdot)\in L^{\infty}(\Omega)$; $\lambda>0$ and $\alpha\in(0,1)$ are real parameters.
Problems involving the double phase operator have been extensively studied in the literature due to their significant applications in various fields. The double phase operator, defined as
$\mathrm{div}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u),\quad u\in W_{0}^{1,\mathcal{H}}(\Omega),$
is intrinsically connected to the two-phase integral functional
$u\to\int_{\Omega}\left(|\nabla u|^{p(x)}+\mu(x)|\nabla u|^{q(x)}\right)\mathrm{d}x,\quad u\in W_{0}^{1,\mathcal{H}}(\Omega),$
where $W_{0}^{1,\mathcal{H}}(\Omega)$ denotes a Musielak-Orlicz Sobolev space. This functional exhibits changing ellipticity properties depending on the set where the weight function $\mu(\cdot)$ vanishes, leading to two distinct phases of elliptic behavior. The pioneering work of Zhikov [36] first investigated this functional with constant exponents to model strongly anisotropic materials. In elasticity theory, $\mu(\cdot)$ encodes geometric information about composites consisting of two different materials with power hardening exponents $p(\cdot)$ and $q(\cdot)$ [37]. Early mathematical treatments of these two-phase integrals were conducted by Baroni-Colombo-Mingione [8, 9, 10], Colombo-Mingione [15, 16], and Ragusa-Tachikawa [32], with further contributions from De Filippis-Mingione [22] for nonautonomous integrals.
While numerous authors have established existence and multiplicity results for double phase problems with constant exponents [11, 13, 23, 24, 27, 29, 30, 35], the variable exponent case remains comparatively less explored. Recent contributions in this direction include works by Amoroso-Bonanno-D’Aguì-Winkert [6], Abergi-Bennouna-Benslimane-Ragusa [1], Bahrouni-Radulescu-Winkert [7], Cen-Kim-Kim-Zeng [12], Crespo-Blanco-Gasinski-Harjulehto-Winkert [17], Leonardi-Papageorgiou [26], Liu-Pucci [28], Kim-Kim-Oh-Zeng [25], Vetro-Winkert [33], Zeng-Radulescu-Winkert [34], and Avci [5].
The distinctive feature of problem ($\mathcal{P}_{\lambda}$) is the presence of the singular Hardy-type potential in the term $\mathcal{B}(x,\mu,u)$, which introduces critical singularities at the origin. This represents a significant departure from the existing literature on double phase problems. Physically, such singular potentials model phenomena involving strongly attractive forces or singular sources, appearing in quantum mechanics, molecular physics, and relativistic field theories. In material science, they can describe composites with point defects or singular material properties that intensify near specific locations. Mathematically, the fractional-type-Hardy potential introduces substantial challenges due to its critical nature and the lack of compactness in the associated energy functional which requires a careful treatment of the underlying Musielak-Orlicz Sobolev spaces.
Our work aims to bridge the gap between double phase operators and singular potentials in the variable exponent setting, establishing new existence results that extend the current understanding of these complex nonlinear phenomena.
2 Mathematical Background and Preliminaries
We start with some basic concepts of variable Lebesgue-Sobolev spaces. For more details, and the proof of the following propositions, we refer the reader to [18, 19, 21, 31].
Define
$C_{+}\left(\overline{\Omega}\right)=\left\{h\in C\left(\overline{\Omega}\right),h(x)>1\text{ for all\ }x\in\overline{\Omega}\right\}.$
For $h\in C_{+}(\overline{\Omega})$ denote
$h^{-}:=\underset{x\in\overline{\Omega}}{\min}h(x),\quad h^{+}:=\underset{x\in\overline{\Omega}}{\max}h(x)<\infty.$
For any $h\in C_{+}\left(\overline{\Omega}\right)$, we define the variable exponent Lebesgue space by
$L^{h(x)}(\Omega)=\left\{u\mid u:\Omega\rightarrow\mathbb{R}\text{ is measurable},\int_{\Omega}|u(x)|^{h(x)}dx<\infty\right\}.$
Then, $L^{h(x)}(\Omega)$ endowed with the norm
$|u|_{h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left|\frac{u(x)}{\lambda}\right|^{h(x)}dx\leq 1\right\},$
becomes a Banach space.
Proposition 2.1 (Hölder Inequality).
For any $u\in L^{h(x)}(\Omega)$ and $v\in L^{h^{\prime}(x)}(\Omega)$, we have
$\int_{\Omega}|uv|dx\leq C(h^{-},(h^{-})^{\prime})|u|_{h(x)}|v|_{h^{\prime}(x)},$
where $L^{h^{\prime}(x)}(\Omega)$ is conjugate space of $L^{h(x)}(\Omega)$ such that $\frac{1}{h(x)}+\frac{1}{h^{\prime}(x)}=1$.
The convex functional $\rho:L^{h(x)}(\Omega)\rightarrow\mathbb{R}$ defined by
$\rho(u)=\int_{\Omega}|u(x)|^{h(x)}dx,$
is called modular on $L^{h(x)}(\Omega)$.
Proposition 2.2.
If $u,u_{n}\in L^{h(x)}(\Omega)$, we have
  • $(i)$
    $|u|_{h(x)}<1(=1;>1)\Leftrightarrow\rho(u)<1(=1;>1);$
  • $(ii)$
    $|u|_{h(x)}>1\implies|u|_{h(x)}^{h^{-}}\leq\rho(u)\leq|u|_{h(x)}^{h^{+}}$;
    $|u|_{h(x)}\leq 1\implies|u|_{h(x)}^{h^{+}}\leq\rho(u)\leq|u|_{h(x)}^{h^{-}};$
  • $(iii)$
    $\lim\limits_{n\rightarrow\infty}|u_{n}|_{h(x)}=0\Leftrightarrow\lim\limits_{n\rightarrow\infty}\rho(u_{n})=0;\newline \lim\limits_{n\rightarrow\infty}|u_{n}|_{h(x)}=\infty\Leftrightarrow\lim\limits_{n\rightarrow\infty}\rho(u_{n})=\infty$.
Proposition 2.3.
If $u,u_{n}\in L^{h(x)}(\Omega)$, then the following statements are equivalent:
  • $(i)$
    $\lim\limits_{n\rightarrow\infty}|u_{n}-u|_{h(x)}=0$;
  • $(ii)$
    $\lim\limits_{n\rightarrow\infty}\rho(u_{n}-u)=0$;
  • $(iii)$
[…]
The variable exponent Sobolev space $W^{1,h(x)}(\Omega)$ is defined by
$W^{1,h(x)}(\Omega)=\{u\in L^{h(x)}(\Omega):|\nabla u|\in L^{h(x)}(\Omega)\},$
with the norm
$\|u\|_{1,h(x)}=|u|_{h(x)}+|\nabla u|_{h(x)},$
or equivalently
$\|u\|_{1,h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left(\left|\frac{\nabla u(x)}{\lambda}\right|^{h(x)}+\left|\frac{u(x)}{\lambda}\right|^{h(x)}\right)dx\leq 1\right\},$
for all $u\in W^{1,h(x)}(\Omega)$.
Proposition 2.4.
If $1<h^{-}\leq h^{+}<\infty$, then the spaces $L^{h(x)}(\Omega)$ and $W^{1,h(x)}(\Omega)$ are separable and reflexive Banach spaces.
The space $W_{0}^{1,h(x)}(\Omega)$ is defined as
$W_{0}^{1,h(x)}(\Omega):=\bigcap_{C_{0}^{\infty}(\Omega)\subset\mathcal{K}}\mathcal{K}\quad\textit{ for closed subsets }\mathcal{K}\subset W^{1,h(x)}(\Omega).$
Therefore, $W_{0}^{1,h(x)}(\Omega)$ is also a separable and reflexive Banach space. Furthermore, since Poincaré inequality holds on $W_{0}^{1,h(x)}(\Omega)$, we have
$|\nabla u|_{h(x)}\leq\|u\|_{1,h(x)}\leq(C+1)|\nabla u|_{h(x)},\quad\forall u\in W_{0}^{1,h(x)}(\Omega),\quad 0<C\in\mathbb{R},$
which implies $\|u\|_{1,h(x)}$ and $|\nabla u|_{h(x)}$ are equivalent norms on $W_{0}^{1,h(x)}(\Omega)$. Therefore, for any $u\in W_{0}^{1,h(x)}(\Omega)$ we can define an equivalent norm $\|u\|$ such that
$\|u\|=|\nabla u|_{h(x)}.$
Proposition 2.5.
Let $r\in C(\overline{\Omega})$. If $1\leq r(x)<h^{\ast}(x)$ for all $x\in\overline{\Omega}$, then the embeddings $W^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ and $W_{0}^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ are compact and continuous, where
$h^{\ast}(x)=\left\{\begin{array}[]{cc}\frac{Nh(x)}{N-h(x)}&\text{if }h(x)<N,\\ +\infty&\text{if }h(x)\geq N.\end{array}\right$.
In the sequel, we introduce the double phase operator, the Musielak–Orlicz space, and the Musielak–Orlicz Sobolev space, respectively.
Throughout the paper, we assume the following.
  • $(H_{1})$
    $p,q\in C(\overline{\Omega})$, $1<p(x),q(x)<N$ with $p(x)<q(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
  • $(H_{2})$
    $\mu\in L^{\infty}(\Omega)$ such that $\mu(\cdot)\geq 0$.
Let $\mathcal{H}:\Omega\times[0,\infty]\to[0,\infty]$ be the nonlinear function defined by
$\mathcal{H}(x,t)=t^{p(x)}+\mu(x)t^{q(x)}\ \text{for all}\ (x,t)\in\Omega\times[0,\infty).$
Then the corresponding modular $\rho_{\mathcal{H}}(\cdot)$ is given by
$\rho_{\mathcal{H}}(u)=\int_{\Omega}\mathcal{H}(x,|u|)dx=\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx.$
The Musielak-Orlicz space $L^{\mathcal{H}}(\Omega)$, is defined by
$L^{\mathcal{H}}(\Omega)=\left\{u:\Omega\to\mathbb{R}\,\,\text{measurable};\,\,\rho_{\mathcal{H}}(u)<+\infty\right\},$
endowed with the Luxemburg norm
$\|u\|_{\mathcal{H}}=\inf\left\{\hat{\lambda}>0:\rho_{\mathcal{H}}\left(\frac{u}{\hat{\lambda}}\right)\leq 1\right\}.$
Analogous to Proposition 2.2, there are similar relationship between the modular $\rho_{\mathcal{H}}(\cdot)$ and the norm $\|\cdot\|_{\mathcal{H}}$, see [17, Proposition 2.13] for a detailed proof. Crespo-Blanco
Proposition 2.6.
Assume $(H_{1})$ hold, $u\in L^{\mathcal{H}}(\Omega)$ and $\hat{\lambda}\in\mathbb{R}$. Then
  • $(i)$
    If $u\neq 0$, then $\|u\|_{\mathcal{H}}=\hat{\lambda}\Leftrightarrow\rho_{\mathcal{H}}(\frac{u}{\hat{\lambda}})=1$,
  • $(ii)$
    $\|u\|_{\mathcal{H}}<1\ (\text{resp.}\ >1,=1)\Leftrightarrow\rho_{\mathcal{H}}(\frac{u}{\hat{\lambda}})<1\ (\text{resp.}\ >1,=1)$,
  • $(iii)$
    If $\|u\|_{\mathcal{H}}<1\Rightarrow\|u\|_{\mathcal{H}}^{q^{+}}\leq\rho_{\mathcal{H}}(u)\leq\|u\|_{\mathcal{H}}^{p^{-}}$,
  • $(iv)$
    If $\|u\|_{\mathcal{H}}>1\Rightarrow\|u\|_{\mathcal{H}}^{p^{-}}\leq\rho_{\mathcal{H}}(u)\leq\|u\|_{\mathcal{H}}^{q^{+}}$,
  • $(v)$
    $\|u\|_{\mathcal{H}}\to 0\Leftrightarrow\rho_{\mathcal{H}}(u)\to 0$,
  • $(vi)$
    $\|u\|_{\mathcal{H}}\to+\infty\Leftrightarrow\rho_{\mathcal{H}}(u)\to+\infty$,
  • $(vii)$
    $\|u\|_{\mathcal{H}}\to 1\Leftrightarrow\rho_{\mathcal{H}}(u)\to 1$,
  • $(viii)$
    If $u_{n}\to u$ in $L^{\mathcal{H}}(\Omega)$, then $\rho_{\mathcal{H}}(u_{n})\to\rho_{\mathcal{H}}(u)$.
The Musielak-Orlicz Sobolev space $W^{1,\mathcal{H}}(\Omega)$ is defined by
$W^{1,\mathcal{H}}(\Omega)=\left\{u\in L^{\mathcal{H}}(\Omega):|\nabla u|\in L^{\mathcal{H}}(\Omega)\right\},$
and equipped with the norm
$\|u\|_{1,\mathcal{H}}=\|\nabla u\|_{\mathcal{H}}+\|u\|_{\mathcal{H}},$
where $\|\nabla u\|_{\mathcal{H}}=\|\,|\nabla u|\,\|_{\mathcal{H}}$, and $|\nabla u|$ is the $n$-dimensional Euclidean norm of the gradient of $u$.
The space $W_{0}^{1,\mathcal{H}}(\Omega)$ is defined (in a similar way as with $W_{0}^{1,h(x)}(\Omega)$) as $W_{0}^{1,\mathcal{H}}(\Omega):=\overline{C_{0}^{\infty}(\Omega)}^{\|\cdot\|_{1,\mathcal{H}}}$. Notice that $L^{\mathcal{H}}(\Omega),W^{1,\mathcal{H}}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)$ are uniformly convex and reflexive Banach spaces. Moreover we have the following embeddings [17, Proposition 2.16].
Proposition 2.7.
Let $(H)$ be satisfied. Then the following embeddings hold:
  • $(i)$
    $L^{\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega),W^{1,\mathcal{H}}(\Omega)\hookrightarrow W^{1,h(x)}(\Omega)$, $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,h(x)}(\Omega)$ are continuous for all $h\in C(\overline{\Omega})$ with $1\leq h(x)\leq p(x)$ for all $x\in\overline{\Omega}$.
  • $(ii)$
    $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ are compact for all $h\in C(\overline{\Omega})$ with $1\leq h(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
As a conclusion of Proposition 2.7, we have the continuous embedding $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$, and $C_{\mathcal{H}}$ denotes the best constant such that
$\|u\|_{h(x)}\leq C_{\mathcal{H}}\|u\|_{1,\mathcal{H},0}.$
Moreover, by [17, Proposition 2.18], $W_{0}^{1,\mathcal{H}}(\Omega)$ is compactly embedded in $L^{\mathcal{H}}(\Omega)$. Thus, the space $W_{0}^{1,\mathcal{H}}(\Omega)$ can be equipped with the equivalent norm
$\|u\|_{1,\mathcal{H},0}=\|\nabla u\|_{\mathcal{H}}.$
Proposition 2.8.
The convex functional
$\varrho_{\mathcal{H}}(u):=\int_{\Omega}\left(\frac{|\nabla u|^{p(x)}}{p(x)}+\mu(x)\frac{|\nabla u|^{q(x)}}{q(x)}\right)dx$
is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$, and its derivative $\varrho^{\prime}_{\mathcal{H}}$ satisfies the ($S_{+}$)-property with the derivative
$\langle\varrho^{\prime}_{\mathcal{H}}(u),\varphi\rangle=\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla\varphi dx,$
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, where $\langle\cdot,\cdot\rangle$ is the dual pairing between $W_{0}^{1,\mathcal{H}}(\Omega)$ and its dual $W_{0}^{1,\mathcal{H}}(\Omega)^{*}$ [17].
Next, we present and prove a fractional-Hardy-type inequality which we use to keep the singular integral $\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$ under the control so that it does not exhibit an uncontrolled growth (blow up). When $\alpha\equiv 1$, the multi-phase case of (1) is given in [3], while the anisotropic case is given in [4].
Lemma 2.9.
Assume that $(H_{1})$ holds. If, additionally, $\alpha p^{-}\neq 1$ and $\alpha q^{+}\neq 1$, then the following hold:
$(i)$ There exists a positive constant $C_{N}(p^{-},q^{+},\alpha)$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega),$ (1)
holds, where $C_{N}(p^{-},q^{+},\alpha):=\frac{(1+\|\mu\|_{\infty})}{p^{-}}\max\{C_{N}(q^{+}),C_{N}(p^{-})\}$ with $C_{N}(q^{+})=\hat{c}_{1}\left(\frac{q^{+}}{N-\alpha q^{+}}\right)^{q^{+}}$, $C_{N}(p^{-})=\hat{c}_{2}\left(\frac{p^{-}}{N-\alpha p^{-}}\right)^{p^{-}}$ where $\hat{c}_{1}>0,\hat{c}_{2}>0$ are the embedding constants not depending on $u$; and $\kappa=p^{-}$ if $\|u\|_{1,\mathcal{H},0}<1$, $\kappa=q^{+}$ if $\|u\|_{1,\mathcal{H},0}\geq 1$.
$(ii)$ There exists $0<M\in\mathbb{R}$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha\tau}}\|u\|^{\kappa}_{\mathcal{H}},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega)$ (2)
holds, where $\tau=p^{-}$ if $0<|x|<M<1$, and $\tau=q^{+}$ if $1\leq|x|<M$.
Proof.
$(i)$ We split $\Omega$ in the following two subsets
$\Omega_{<}:=\{x\in\Omega:|x|<1\},\quad\Omega_{\geq}:=\{x\in\Omega:|x|\geq 1\}.$ (3)
Since $\Omega$ is bounded and connected, we obviously have $\Omega=\Omega_{<}\cup\Omega_{\geq}$.
Case I: $x\in\Omega_{<}$.
In this case, we have $|x|^{\alpha}>|x|$. Thus, we need to consider the following orderings:
  • (a)
    $|u(x)|\geq|x|^{\alpha}>|x|$,
  • (b)
    $|x|^{\alpha}>|u(x)|>|x|$,
  • (c)
    $|x|^{\alpha}>|x|\geq|u(x)|$.
For (a): $|u(x)|\geq|x|^{\alpha}$. Using $(H_{1})$, for all $x\in\Omega_{<}$, $x\neq 0$, we have
$\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\leq\frac{1}{p^{-}}\left(\frac{|u(x)|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u(x)|^{q(x)}}{|x|^{\alpha q(x)}}\right)\leq\frac{\left(1+\|\mu\|_{\infty}\right)}{p^{-}}\frac{|u(x)|^{q^{+}}}{|x|^{\alpha q^{+}}}.$ (4)
Henceforth, we follow the same reasoning as with [3].
We start with the identity (adapted from Theorem 7, [20])
$\frac{|u(x)|^{q^{+}}}{|x|^{\alpha q^{+}}}=\frac{1}{(1-\alpha q^{+})}|u(x)|^{q^{+}}\frac{x}{|x|}\cdot\nabla\left[\frac{1}{|x|^{\alpha q^{+}-1}}\right].$ (5)
Using the divergence theorem gives
$\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx=\frac{1}{(\alpha q^{+}-1)}\int_{\Omega_{<}}\nabla\cdot\left[|u|^{q^{+}}\frac{x}{|x|}\right]\frac{1}{|x|^{\alpha q^{+}-1}}dx+\frac{1}{(1-\alpha q^{+})}\int_{\partial\Omega}|u|^{q^{+}}\frac{x}{|x|^{\alpha q^{+}}}\cdot\textbf{n}dS,$ (6)
which is equal to
$(N-\alpha q^{+})\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx=(-q^{+})\int_{\Omega_{<}}|u|^{q^{+}-2}u\nabla u\cdot\frac{x}{|x|^{\alpha q^{+}}}dx.$ (7)
By Cauchy-Schwarz inequality, we have
$(N-\alpha q^{+})\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\leq q^{+}\int_{\Omega_{<}}\frac{|u|^{q^{+}-1}}{|x|^{\alpha q^{+}-1}}|\nabla u|dx.$ (8)
Then the Hölder inequality provides
$\int_{\Omega_{<}}\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\leq\left(\frac{q^{+}}{N-\alpha q^{+}}\right)^{q^{+}}\int_{\Omega_{<}}|\nabla u|^{q^{+}}dx.$ (9)
By the continuous embedding $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,q^{+}}(\Omega)$, we get
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(q^{+})\|u\|_{1,\mathcal{H},0}^{q^{+}}.$ (10)
For (b) and (c): $|x|^{\alpha}>|u(x)|$. Considering $(H_{1})$ again, for all $x\in\Omega_{<}$, $x\neq 0$, it reads
$\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\leq\frac{(1+\|\mu\|_{\infty})}{p^{-}}\frac{|u(x)|^{p^{-}}}{|x|^{\alpha p^{-}}}.$ (11)
Applying the same reasoning and steps of (a), it follows
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-})\|u\|_{1,\mathcal{H},0}^{p^{-}}.$ (12)
As a result of (10) and (12), it holds
$\int_{\Omega_{<}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega_{<}),$ (13)
Case II: $x\in\Omega_{\geq}$.
In this case, we have $|x|\geq|x|^{\alpha}$, and hence we consider the following:
  • (a’)
    $|u(x)|\geq|x|\geq|x|^{\alpha}$,
  • (b’)
    $|x|>|u(x)|\geq|x|^{\alpha}$,
  • (c’)
    $|x|\geq|x|^{\alpha}>|u(x)|$.
Then for (a’) and (b’), we have $\frac{|u(x)|}{|x|^{\alpha}}\geq 1$; whereas, for (c’), we have $\frac{|u(x)|}{|x|^{\alpha}}<1$. Both cases follow in the same way as with Case I, so we can omit the details and write
$\int_{\Omega_{\geq}}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq C_{N}(p^{-},q^{+},\alpha)\|u\|_{1,\mathcal{H},0}^{\kappa},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega_{\geq}),$ (14)
By combining (13) and (14), we can conclude that (1) holds for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$.
$(ii)$ There exists $M>0$ such that $0<|x|<M$ for all $x\in\Omega\setminus\{0\}$. Thus, $\max\{|x|^{\alpha p(x)},|x|^{\alpha q(x)}\}<M^{\alpha\tau}$ which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha\tau}}\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx>\frac{1}{M^{\alpha\tau}}\|u\|^{\kappa}_{\mathcal{H}}.$ (15)
In the sequel, we let $\mathcal{G}(u):=\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$.
Definition 2.10.
Let $X$ be a Banach space, and $J:X\rightarrow\mathbb{R}$ be a $C^{1}$-functional. We say that $J$ satisfies the Palais-Smale condition ($(\mathbf{PS})$ for short) if: every sequence $(u_{n})\subset X$ such that $J(u_{n})$ is bounded and $J^{\prime}(u_{n})\rightarrow 0$ in $X^{*}$ admits a convergent subsequence in $X$.
3 Variational Framework and Main Results
We define the energy functional $\mathcal{I}:W_{0}^{1,\mathcal{H}}(\Omega)\rightarrow\mathbb{R}$ corresponding to equation ($\mathcal{P}_{\lambda}$) by
$\mathcal{I}(u)=$ $\int_{\Omega}\left(\frac{|\nabla u|^{p(x)}}{p(x)}+\mu(x)\frac{|\nabla u|^{q(x)}}{q(x)}\right)dx+\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$-\lambda\int_{\Omega}F(x,u)dx,$
or
$\mathcal{I}(u)=\varrho_{\mathcal{H}}(u)+\mathcal{G}(u)-\lambda\int_{\Omega}F(x,u)dx.$
Definition 3.1.
A function $u\in W_{0}^{1,\mathcal{H}}(\Omega)$ is called a (weak) solution to problem ($\mathcal{P}_{\lambda}$) if
$\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla\varphi dx+\int_{\Omega}\left(\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}\right)\varphi dx$
$=\lambda\int_{\Omega}f(x,u)\varphi dx,\,\,\ \forall\varphi\in W_{0}^{1,\mathcal{H}}(\Omega),$ (16)
where $F(x,t)=\int_{0}^{t}f(x,s)ds$.
Lemma 3.2.
$\mathcal{I}\in C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$ with the derivative
$\langle\mathcal{I}^{\prime}(u),\varphi\rangle$ $=\langle\varrho^{\prime}_{\mathcal{H}}(u),\varphi\rangle+\langle\mathcal{G}^{\prime}(u),\varphi\rangle-\lambda\langle f(x,u),\varphi\rangle,\,\,\ \forall\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$ (17)
Moreover, the critical points of $\mathcal{I}$ are the solutions of problem ($\mathcal{P}_{\lambda}$).
Proof.
By Proposition 2.8, $\varrho_{\mathcal{H}}$ is a continuously Gâteaux differentiable functional with the derivative $\langle\varrho^{\prime}_{\mathcal{H}}(\cdot),\varphi\rangle$. Moreover, using the assumption $(\mathbf{f}_{1})$, and applying Holder inequality along with the related embeddings shows that the functional $\int_{\Omega}F(x,u)dx$ is also continuously Gâteaux differentiable with the derivative $\langle f(x,u),\varphi\rangle=\int_{\Omega}f(x,u)\varphi dx$. Thus, it remains only to determine the Gâteaux derivative of $\mathcal{G}$ and show that it has the same regularity on $W_{0}^{1,\mathcal{H}}(\Omega)$.
To begin, by the mean value theorem, there are $0<\varepsilon_{1},\varepsilon_{2}<1$ such that
$\langle\mathcal{G}^{\prime}(u),\varphi\rangle$ $=\lim_{t\to 0}\int_{\Omega}\frac{1}{t}\left(\frac{|u+t\varphi|^{p(x)}-|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{\mu(x)|u+t\varphi|^{q(x)}-|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$=\lim_{t\to 0}\int_{\Omega}\frac{1}{t}\left(\frac{d}{d\gamma}\frac{|u+\gamma t\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\bigg|_{\gamma=\varepsilon_{1}}+\frac{d}{d\gamma}\frac{\mu(x)|u+\gamma t\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\bigg|_{\gamma=\varepsilon_{2}}\right)dx$
$=\lim_{t\to 0}\int_{\Omega}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\right$.
$+\left.|x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\right)\varphi dx,$
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, $\gamma\in\mathbb{R}$. Using Young inequality, we obtain
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|$ $\leq k|x|^{-\alpha p(x)}(|u|^{p(x)}+|\varphi|^{p(x)}),$ (19)
and
$\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$ $\leq k|x|^{-\alpha q(x)}(\mu(x)|u|^{q(x)}+\mu(x)|\varphi|^{q(x)}).$ (20)
where $k:=\frac{2^{q^{+}}(q^{+}-1)}{p^{-}}$. Using (19),(20) and Lemma 2.9, we can write
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|+\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$
$\leq k(|x|^{-\alpha p(x)}|u|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})+k(|x|^{-\alpha p(x)}|\varphi|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})$
$\leq kq^{+}C_{N}(p^{-},q^{+},\alpha)\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (21)
Therefore, by the Lebesgue dominated convergence theorem it reads
$\langle\mathcal{G}^{\prime}(u),\varphi\rangle$ $=\int_{\Omega}\lim_{t\to 0}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\right$.
$+\left.|x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\right)\varphi dx,$
$=\int_{\Omega}\left(|x|^{-\alpha p(x)}|u|^{p(x)-2}u+\mu(x)|x|^{-\alpha q(x)}|u|^{q(x)-2}u\right)\varphi dx.$ (22)
Since the right-hand side of (3), as a function of $\varphi$, is a linear functional, $\mathcal{G}^{\prime}$ is linear on $W_{0}^{1,\mathcal{H}}(\Omega)$.
Next, by Young inequality and Lemma 2.9, we obtain
$|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|$ $\leq\int_{\Omega}\left(|x|^{-p(x)}|u|^{p(x)-1}|\varphi|+|x|^{-\alpha q(x)}\mu(x)|u|^{q(x)-1}|\varphi|\right)dx$
$\leq(q^{+}-1)\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{|\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\right)+\left(\frac{\mu(x)|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}+\frac{\mu(x)|\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$\leq(q^{+}-1)C_{N}(p^{-},q^{+},\alpha)\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (23)
Therefore, for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$, we have
$\|\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|\leq(q^{+}-1)C_{N}(p^{-},q^{+},\alpha)(\|u\|_{1,\mathcal{H},0}^{\kappa}+1),$ (24)
which means that $\mathcal{G}^{\prime}$ is bounded. Therefore, $\mathcal{G}$ is Gâteaux differentiable whose derivative is given by the formula (3). Now, we proceed with the continuity of $\mathcal{G}^{\prime}$. Let $(u_{n})\subset W_{0}^{1,\mathcal{H}}(\Omega)$ such that $u_{n}\to u$ in $W_{0}^{1,\mathcal{H}}(\Omega)$. Applying the inequality (Proposition 17.2, [14])
$\left|\left|\xi\right|^{s-2}\xi-\left|\psi\right|^{s-2}\psi\right|\leq C_{s}\left|\xi-\psi\right|\{|\xi|+|\psi|\}^{s-2},\quad\xi,\psi\in\mathbb{R}^{N},\,\,s>1,$ (25)
gives
$\left|\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u),\varphi\rangle\right|$ $\leq\int_{\Omega}|x|^{-p(x)}\left||u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right||\varphi|dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left||u_{n}|^{q(x)-2}u_{n}-|u|^{q(x)-2}u\right||\varphi|dx$
$\leq C_{p}\int_{\Omega}|x|^{-p(x)}\{|u_{n}|+|u|\}^{p(x)-2}|u_{n}-u||\varphi|dx$
$+C_{q}\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\{|u_{n}|+|u|\}^{q(x)-2}|u_{n}-u||\varphi|dx.$ (26)
Thus, considering that $u_{n}\to u$ in $L(\Omega)$ and $(u_{n})$ is bounded, it reads
$\|\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}\left|\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u),\varphi\rangle\right|\to 0.$ (27)
Therefore, $\mathcal{G}$ is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$.
Finally, $\mathcal{I}^{\prime}(u)=0$ is exactly the weak formulation of ($\mathcal{P}_{\lambda}$), so critical points of $\mathcal{I}$ are weak solutions. This completes the proof. ∎
The main results of the present paper is given below.
Theorem 3.3.
Assume that the following assumptions hold:
  • $(\mathbf{\beta_{0}})$
    $\beta\in C_{+}\left(\overline{\Omega}\right)$ such that $q^{+}<\beta^{-}\leq\beta^{+}<p^{\ast}(x)$   $\forall x\in\overline{\Omega}$.
  • $(\mathbf{f}_{1})$
    $f:\overline{\Omega}\times\mathbb{R}\rightarrow\mathbb{R}$ is a Carathéodory function and satisfies the growth condition
    $\left|f(x,t)\right|\leq c_{1}+c_{2}\left|t\right|^{\beta(x)-1},\quad\forall\left(x,t\right)\in\overline{\Omega}\times\mathbb{R},$
    where $c_{1}$ and $c_{2}$ are positive constants.
  • $(\mathbf{f}_{2})$
    $f(x,t)=o\left(\left|t\right|^{q^{+}-1}\right)$, $t\rightarrow 0$ uniformly $\forall x\in\overline{\Omega}$.
  • $(\mathbf{AR})$
    $\exists K>0$, $\theta>q^{+}$ such that
    $0<\theta F(x,t)\leq f(x,t)t,\quad\left|t\right|\geq K\text{ a.e.}\,x\in\overline{\Omega}.$
Then problem ($\mathcal{P}_{\lambda}$) has at least one nontrivial positive weak solution.
To obtain the result of Theorem 3.3, we need to show Lemma 3.4 and Lemma 3.5 hold.
Lemma 3.4.
Suppose $(\mathbf{\beta_{0}})$, $(\mathbf{f}_{1})$, $(\mathbf{f}_{2})$ and $(\mathbf{AR})$ hold. Then the following statements hold:
  • $(i)$
    There exist two positive real numbers $\gamma$ and $\eta$ such that $\mathcal{I}(u)\geq\eta>0$, for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$ with $\|u\|_{1,\mathcal{H},0}=\gamma$.
  • $(ii)$
    There exists a $\hat{u}\in W_{0}^{1,\mathcal{H}}(\Omega)$ such that $\|\hat{u}\|_{1,\mathcal{H},0}>\gamma$  and   $\mathcal{I}(\hat{u})<0$.
Proof.
$(i)$ By $(\mathbf{f}_{1})$ and $(\mathbf{f}_{2})$, one can write
$F(x,t)\leq\varepsilon|t|^{q^{+}}+C_{\varepsilon}|t|^{\beta(x)},\quad\forall\left(x,t\right)\in\overline{\Omega}\times\mathbb{R}.$ (28)
Then
$\mathcal{I}(u)$ $\geq\frac{1}{q^{+}}\rho_{\mathcal{H}}(\nabla u)+\frac{1}{q^{+}M^{\alpha\tau}}\rho_{\mathcal{H}}(u)-\lambda\varepsilon\int_{\Omega}|u|^{q^{+}}dx-\lambda C_{\varepsilon}\int_{\Omega}|u|^{\beta(x)}dx$
$\geq\frac{1}{q^{+}}\rho_{\mathcal{H}}(\nabla u)-\lambda\varepsilon|u|^{q^{+}}_{q^{+}}-\lambda C_{\varepsilon}\left(|u|^{\beta^{+}}_{\beta^{+}}+|u|^{\beta^{-}}_{\beta^{-}}\right).$ (29)
Using Proposition 2.7, and assuming that $\|u\|_{1,\mathcal{H},0}<1$ provides
$\mathcal{I}(u)$ $\geq\frac{1}{q^{+}}\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda\varepsilon c_{3}\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda C_{\varepsilon}c_{4}\|u\|_{1,\mathcal{H},0}^{\beta^{-}}$
$=\left(\frac{1}{q^{+}}-\lambda\varepsilon c_{3}\right)\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda C_{\varepsilon}c_{4}\|u\|_{1,\mathcal{H},0}^{\beta^{-}}.$ (30)
[…]
$(ii)$ Due to $(\mathbf{AR})$, we can write
$F(x,t)\geq c_{5}|t|^{\theta},\quad\forall x\in\overline{\Omega},\,\,|t|\geq K.$ (31)
Thus, for $0\neq\phi\in W_{0}^{1,\mathcal{H}}(\Omega)$ and $t>1$, we have
$\mathcal{I}(t\phi)$ $\leq\frac{t^{q^{+}}}{p^{-}}\left(\rho_{\mathcal{H}}(\nabla\phi)+\int_{\Omega}\left(\frac{|\phi|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|\phi|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\right)-\lambda c_{5}t^{\theta}\int_{\Omega}|u|^{\theta}dx,$ (32)
which implies that $\mathcal{I}(t\phi)\to$-$\infty$ as $t\to\infty$. Therefore, letting $\hat{u}=t\phi$, and hence $\|\hat{u}\|_{1,\mathcal{H},0}=t\|\phi\|_{1,\mathcal{H},0}$, leads to $\|\hat{u}\|_{1,\mathcal{H},0}>1>\gamma$ and $\mathcal{I}(\hat{u})<0$ provided $t$ is large enough,. ∎
Lemma 3.5.
Suppose $(\mathbf{\beta_{0}})$, $(\mathbf{f}_{1})$ and $(\mathbf{AR})$ hold. Then $\mathcal{I}$ satisfies the $(\mathbf{PS})$ condition.
Proof.
Let assume that there exists a sequence $(u_{n})\subset W_{0}^{1,\mathcal{H}}(\Omega)$ such that
$|\mathcal{I}(u_{n})|\leq c_{0}\quad\text{ and }\quad\|\mathcal{I}^{\prime}(u_{n})\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}\rightarrow 0,\quad 0<c_{0}\in\mathbb{R}.$ (33)
Using (33), $(\mathbf{AR})$, and Proposition 2.6, we obtain
$c_{0}$ $\geq\mathcal{I}(u_{n})=\varrho_{\mathcal{H}}(u_{n})+\mathcal{G}(u_{n})-\lambda\int_{\Omega}F(x,u_{n})dx$
$\geq\left(\frac{1}{q^{+}}-\frac{1}{\theta}\right)\rho_{\mathcal{H}}(\nabla u_{n})+\frac{1}{\theta}\int_{\Omega}\left(\rho_{\mathcal{H}}(\nabla u_{n})+\frac{|u_{n}|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u_{n}|^{q(x)}}{|x|^{\alpha q(x)}}-\lambda f(x,u_{n})u_{n}\right)dx$
$\geq\left(\frac{1}{q^{+}}-\frac{1}{\theta}\right)\|u_{n}\|_{1,\mathcal{H},0}^{p^{-}}-\frac{1}{\theta}\|u_{n}\|_{1,\mathcal{H},0}\|\mathcal{I}^{\prime}(u_{n})\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}.$ (34)
Therefore $(u_{n})$ is bounded in $W_{0}^{1,\mathcal{H}}(\Omega)$. Since $W_{0}^{1,\mathcal{H}}(\Omega)$ is reflexive, passing to a subsequence, we have $u_{n}\rightharpoonup u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$. Then using $(\mathbf{f}_{1})$, Hölder inequality, and Proposition 2.7, we obtain
$\left|\langle f(x,u_{n}),u_{n}-u_{0}\rangle\right|\leq c_{1}\left||u_{n}|^{\beta(x)-1}\right|_{\frac{\beta(x)}{\beta(x)-1}}\left|u_{n}-u_{0}\right|_{\beta(x)}+c_{2}\int_{\Omega}\left|u_{n}-u_{0}\right|dx\to 0.$ (35)
First let’s write
$\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle=\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle+\langle\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle.$ (36)
Taking into consideration that $\mathcal{G}^{\prime}(u_{0})\in W^{1,\mathcal{H}}(\Omega)^{*}$, we obtain
$\langle\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle=\int_{\Omega}\mathcal{G}^{\prime}(u_{0})(u_{n}-u_{0})dx\to 0.$ (37)
Moreover, since $u_{n}\to u_{0}$ in $L^{2}(\Omega)$, $(u_{n})$ is bounded. Thus
$\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle$ $=\int_{\Omega}|x|^{-\alpha p(x)}\left(|u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right)(u_{n}-u_{0})dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left(|u_{n}|^{q(x)-2}u_{n}-|u_{0}|^{q(x)-2}u_{0}\right)(u_{n}-u_{0})dx$
$\leq\int_{\Omega}|x|^{-\alpha p(x)}\left||u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right||u_{n}-u_{0}|dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left||u_{n}|^{q(x)-2}u_{n}-|u_{0}|^{q(x)-2}u_{0}\right||u_{n}-u_{0}|dx$
$\leq C_{p}\int_{\Omega}|x|^{-\alpha p(x)}\{|u_{n}|+|u_{0}|\}^{p(x)-2}|u_{n}-u_{0}|^{2}dx$
$+C_{q}\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\{|u_{n}|+|u_{0}|\}^{q(x)-2}|u_{n}-u_{0}|^{2}dx\to 0$ (38)
where we applied the inequality (25). Thus
$\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle\to 0.$ (39)
Now, using (33) gives
$\langle\mathcal{I}^{\prime}(u_{n}),u_{n}-u_{0}\rangle$ $=\langle\varrho^{\prime}_{\mathcal{H}}(u_{n}),u_{n}-u_{0}\rangle+\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle-\lambda\langle f(x,u_{n}),u_{n}-u_{0}\rangle\to 0.$ (40)
However, (35), (39) and (40) together implies that
$\langle\varrho^{\prime}_{\mathcal{H}}(u_{n}),u_{n}-u_{0}\rangle\to 0.$ (41)
Thus, by Proposition 2.8 $($$S_{+}$)-property), $u_{n}\to u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$. In conclusion, $\mathcal{I}$ satisfies the $(\mathbf{PS})$ condition. ∎
Proof of Theorem 3.3.
By Lemma 3.4, $\mathcal{I}$ has Mountain-Pass geometry; and by Lemma 3.5, $\mathcal{I}$ satisfies $(\mathbf{PS})$. Also, $\mathcal{I}(0)=0$. Thus, by the Mountain-Pass Theorem [2]:
  • $u_{0}$ is a critical point of $\mathcal{I}$ at mountain pass level $c_{MP}$:
    $c_{MP}:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}\mathcal{I}(\gamma(t))\geq\eta>0,$
    where $\Gamma=\{\gamma\in C([0,1],W^{1,\mathcal{H}}_{0}(\Omega)):\gamma(0)=0,\gamma(1)=\hat{u}\}$.
  • Since $\mathcal{I}(u_{0})\geq\eta>0$, $u_{0}\neq 0$ is a nontrivial critical point of $\mathcal{I}$ at level $c_{MP}$;
    $\mathcal{I}(u_{0})=c_{MP}$ and $\mathcal{I}^{\prime}(u_{0})=0$.
Therefore, by Lemma 3.2, $u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$ corresponds to a nontrivial weak solution of ($\mathcal{P}_{\lambda}$).
Define $f_{+}:\overline{\Omega}\times\mathbb{R}\rightarrow[0,\infty)$ by
$f_{+}(x,t)=\left\{\begin{array}[]{cc}f(x,t)&\text{if }\,t\geq 0,\\ 0&\text{if }\,t<0.\end{array}\right$.
Then clearly the modified energy functional $\mathcal{I_{+}}$ is also of class $C^{1}$ on $W_{0}^{1,\mathcal{H}}(\Omega)$.
Step 1.
Define $u^{-}:=\max\{-u,0\}$. We test (3.1) by replacing $u^{-}$ with $\varphi$ which gives
$\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla u^{-}dx+\int_{\Omega}\left(\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}\right)u^{-}dx$
$=\lambda\int_{\Omega}f(x,u)u^{-}dx$ (42)
for a nontrivial weak solution $u$ of ($\mathcal{P}_{\lambda}$). Then considering the decomposition $\Omega=\{x\in\Omega:u(x)\geq 0\}\cup\{x\in\Omega:u(x)<0\}$ for (3), we have
$0\geq-\int_{\{u<0\}}(|\nabla u|^{p(x)}+\mu(x)|\nabla u|^{q(x)})dx-\int_{\{u<0\}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$
$=\lambda\int_{\{u\geq 0\}}f(x,u)u^{-}dx\geq 0,$ (43)
which implies that $u^{-}\equiv 0$. So, since $u=0$ on $\partial\Omega$, by the weak minimum principle there is a set of positive measure where $u\geq 0$, and $u\not\equiv 0$.
Step 2.
We claim that $u(x)>0$ for all $x\in\Omega$. Suppose, on the contrary, that there exists an interior point $x_{0}\in\Omega$ such that $u(x_{0})=0$. Since $u\geq 0$ a.e., this must be a global minimum of $u$. This, by the strong minimum principle, implies that $u$ is constant. That is, $u\equiv 0$ in a neighborhood of $x_{0}$. However, this contradicts $u\not\equiv 0$. Therefore $u(x)>0$ for all $x\in\Omega$.
Funding
This work was supported by Athabasca University Research Incentive Account [140111 RIA].
ORCID
https://orcid.org/0000-0002-6001-627X