### v3

# Positive weak solutions of a double-phase variable exponent problem with a fractional-Hardy-type singular potential and superlinear nonlinearity

Abstract
In the present paper, we study a double-phase variable exponent problem which is set up within a variational framework including a singular potential of fractional-Hardy-type. We employ the Mountain-Pass theorem and the strong minimum principle to obtain the existence of at least one nontrivial positive weak solution.
keywords:
Positive weak solution; double phase operator; variable exponents; fractional-Hardy-type singularity; the Mountain-Pass theorem; the Ambrosetti–Rabinowitz condition; strong minimum principle.
articletype: Research Paper
1 Introduction
In this article, we study the following singular double-phase variable exponent problem
$\begin{cases}\begin{array}[]{rlll}-\mathcal{D}(x,\mu,\nabla u)+\mathcal{S}(x,\mu,u)&=\lambda f(x,u)\text{ in }\Omega,\\ u&=0\text{ on }\partial\Omega,\end{array}\end{cases}$ ($\mathcal{P}_{\lambda}$)
where
$\mathcal{D}(x,\mu,\nabla u):=\mathrm{div}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u),$
$\mathcal{S}(x,\mu,u):=\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}.$
We assume that $\Omega$ is a bounded domain in $\mathbb{R}^{N}$ $(N\geq 3)$ with Lipschitz boundary $\partial\Omega$; $f$ is a Caratheodry function; $1<p,q\in C(\overline{\Omega})$; $0\leq\mu(\cdot)\in L^{\infty}(\Omega)$; $\lambda>0$ and $\alpha\in(0,1)$ are real parameters.
Problems involving the double phase operator have been extensively studied in the literature due to their significant applications in various fields. The double phase operator, defined as
$\mathrm{div}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u),\quad u\in W_{0}^{1,\mathcal{H}}(\Omega),$
is intrinsically connected to the two-phase integral functional
$u\to\int_{\Omega}\left(|\nabla u|^{p(x)}+\mu(x)|\nabla u|^{q(x)}\right)\mathrm{d}x,\quad u\in W_{0}^{1,\mathcal{H}}(\Omega),$
where $W_{0}^{1,\mathcal{H}}(\Omega)$ denotes a Musielak-Orlicz Sobolev space. This functional exhibits changing ellipticity properties depending on the set where the weight function $\mu(\cdot)$ vanishes, leading to two distinct phases of elliptic behavior. The pioneering work of Zhikov [38] first investigated this functional with constant exponents to model strongly anisotropic materials. In elasticity theory, $\mu(\cdot)$ encodes geometric information about composites consisting of two different materials with power hardening exponents $p(\cdot)$ and $q(\cdot)$ [39]. Early mathematical treatments of these two-phase integrals were conducted by Baroni-Colombo-Mingione [8, 9, 10], Colombo-Mingione [15, 16], and Ragusa-Tachikawa [32], with further contributions from De Filippis-Mingione [22] for nonautonomous integrals.
While numerous authors have established existence and multiplicity results for double phase problems with constant exponents [11, 13, 23, 24, 27, 29, 30, 37], the variable exponent case remains comparatively less explored. Recent contributions in this direction include works by Amoroso-Bonanno-D’Aguì-Winkert [6], Abergi-Bennouna-Benslimane-Ragusa [1], Bahrouni-Radulescu-Winkert [7], Cen-Kim-Kim-Zeng [12], Crespo-Blanco-Gasinski-Harjulehto-Winkert [17], Leonardi-Papageorgiou [26], Liu-Pucci [28], Kim-Kim-Oh-Zeng [25], Vetro-Winkert [33], Zeng-Radulescu-Winkert [36], and Avci [5].
The distinctive feature of problem ($\mathcal{P}_{\lambda}$) is the presence of the singular Hardy-type potential in the term $\mathcal{B}(x,\mu,u)$, which introduces critical singularities at the origin. This represents a significant departure from the existing literature on double phase problems. Physically, such singular potentials model phenomena involving strongly attractive forces or singular sources, appearing in quantum mechanics, molecular physics, and relativistic field theories. In material science, they can describe composites with point defects or singular material properties that intensify near specific locations. Mathematically, the fractional-type-Hardy potential introduces substantial challenges due to its critical nature and the lack of compactness in the associated energy functional which requires a careful treatment of the underlying Musielak-Orlicz Sobolev spaces.
Our work aims to bridge the gap between double phase operators and singular potentials in the variable exponent setting, establishing new existence results that extend the current understanding of these complex nonlinear phenomena.
2 Mathematical Background and Preliminaries
We start with some basic concepts of variable Lebesgue-Sobolev spaces. For more details, and the proof of the following propositions, we refer the reader to [18, 19, 21, 31].
Define
$C_{+}\left(\overline{\Omega}\right)=\left\{h\in C\left(\overline{\Omega}\right):1<h(x)\text{ for all\ }x\in\overline{\Omega}\right\}.$
For $h\in C_{+}(\overline{\Omega})$ denote
$h^{-}:=\underset{x\in\overline{\Omega}}{\min}h(x),\quad h^{+}:=\underset{x\in\overline{\Omega}}{\max}h(x)<\infty.$
For a given $h\in C_{+}\left(\overline{\Omega}\right)$, we define the variable exponent Lebesgue space by
$L^{h(x)}(\Omega)=\left\{u\mid u:\Omega\rightarrow\mathbb{R}\text{ is measurable},\int_{\Omega}|u(x)|^{h(x)}dx<\infty\right\},$
equipped with the Luxemburg norm given by
$|u|_{h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left|\frac{u(x)}{\lambda}\right|^{h(x)}dx\leq 1\right\}.$
With this norm, $(L^{h(x)}(\Omega),|\cdot|_{h(x)})$ becomes a separable and reflexive Banach space.
Proposition 2.1 (Hölder’s Inequality).
For any $u\in L^{h(x)}(\Omega)$ and $v\in L^{h^{\prime}(x)}(\Omega)$, we have
$\int_{\Omega}|uv|dx\leq C(h^{-},(h^{-})^{\prime})|u|_{h(x)}|v|_{h^{\prime}(x)},$
where $L^{h^{\prime}(x)}(\Omega)$ is conjugate space of $L^{h(x)}(\Omega)$ such that $\frac{1}{h(x)}+\frac{1}{h^{\prime}(x)}=1$.
The convex functional $\rho:L^{h(x)}(\Omega)\rightarrow\mathbb{R}$ defined by
$\rho(u)=\int_{\Omega}|u(x)|^{h(x)}dx,$
is called modular on $L^{h(x)}(\Omega)$.
Proposition 2.2.
If $u,u_{n}\in L^{h(x)}(\Omega)$, we have
  • $(i)$
    $|u|_{h(x)}<1(=1;>1)\Leftrightarrow\rho(u)<1(=1;>1);$
  • $(ii)$
    $|u|_{h(x)}>1\implies|u|_{h(x)}^{h^{-}}\leq\rho(u)\leq|u|_{h(x)}^{h^{+}}$;
    $|u|_{h(x)}\leq 1\implies|u|_{h(x)}^{h^{+}}\leq\rho(u)\leq|u|_{h(x)}^{h^{-}};$
  • $(iii)$
    $\lim\limits_{n\rightarrow\infty}|u_{n}|_{h(x)}=0\Leftrightarrow\lim\limits_{n\rightarrow\infty}\rho(u_{n})=0;\newline \lim\limits_{n\rightarrow\infty}|u_{n}|_{h(x)}=\infty\Leftrightarrow\lim\limits_{n\rightarrow\infty}\rho(u_{n})=\infty$.
Proposition 2.3.
If $u,u_{n}\in L^{h(x)}(\Omega)$, then the following statements are equivalent:
  • $(i)$
    $\lim\limits_{n\rightarrow\infty}|u_{n}-u|_{h(x)}=0$;
  • $(ii)$
    $\lim\limits_{n\rightarrow\infty}\rho(u_{n}-u)=0$;
  • $(iii)$
[…]
The variable exponent Sobolev space $W^{1,h(x)}(\Omega)$ is defined by
$W^{1,h(x)}(\Omega)=\{u\in L^{h(x)}(\Omega):|\nabla u|\in L^{h(x)}(\Omega)\},$
with the norm
$\|u\|_{1,h(x)}=|u|_{h(x)}+|\nabla u|_{h(x)},$
or equivalently
$\|u\|_{1,h(x)}=\inf\left\{\lambda>0:\int_{\Omega}\left(\left|\frac{\nabla u(x)}{\lambda}\right|^{h(x)}+\left|\frac{u(x)}{\lambda}\right|^{h(x)}\right)dx\leq 1\right\},$
[…]
The space $W_{0}^{1,h(x)}(\Omega)$ is defined as the closure of $C_{0}^{\infty}(\Omega)$ in $W^{1,h(x)}(\Omega)$.
Proposition 2.4.
If $h\in C_{+}(\overline{\Omega})$ and $h^{+}<\infty$, then the spaces $L^{h(x)}(\Omega)$, $W^{1,h(x)}(\Omega)$, and $W_{0}^{1,h(x)}(\Omega)$ are separable and reflexive Banach spaces.
Furthermore, Poincaré inequality holds in $W_{0}^{1,h(x)}(\Omega)$ [21]; that is, there exists a positive constant $c$ independent of $u$ such that
$|u|_{h(x)}\leq c|\nabla u|_{h(x)},\quad\forall u\in W_{0}^{1,h(x)}(\Omega),$
which implies that $|\nabla u|_{h(x)}$ is an equivalent norm in $W_{0}^{1,h(x)}(\Omega)$. Therefore, on $W_{0}^{1,h(x)}(\Omega)$ we can define an equivalent norm $\|\cdot\|_{h(x)}$ such that
$\|u\|_{h(x)}=|\nabla u|_{h(x)}.$
Proposition 2.5.
If $h\in C_{+}(\overline{\Omega})$, $r\in C(\overline{\Omega})$ and $1\leq r(x)<h^{\ast}(x)$ for all $x\in\overline{\Omega}$, then the embeddings $W^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ and $W_{0}^{1,h(x)}(\Omega)\hookrightarrow L^{r(x)}(\Omega)$ are compact and continuous, where
$h^{\ast}(x)=\left\{\begin{array}[]{cc}\frac{Nh(x)}{N-h(x)}&\text{if }h(x)<N,\\ +\infty&\text{if }h(x)\geq N.\end{array}\right$.
In the sequel, we introduce the double phase operator, the Musielak–Orlicz space, and the Musielak–Orlicz Sobolev space, respectively.
Throughout the paper, we assume the following.
  • $(H_{1})$
    $p,q\in C_{+}(\overline{\Omega})$, $1<p(x)<N$ and $p(x)<q(x)$ for all $x\in\overline{\Omega}$ with
    $q^{+}<\min\{p^{*}(x),N\}$.
  • $(H_{2})$
    $\mu\in L^{\infty}(\Omega)$ such that $\mu(\cdot)\geq 0$, and $\mu(\cdot)\not\equiv 0$.
Let $\mathcal{H}:\Omega\times[0,\infty]\to[0,\infty]$ be the nonlinear function defined by
$\mathcal{H}(x,t)=t^{p(x)}+\mu(x)t^{q(x)}\ \text{for all}\ (x,t)\in\Omega\times[0,\infty).$
Then the corresponding modular $\rho_{\mathcal{H}}(\cdot)$ is given by
$\rho_{\mathcal{H}}(u)=\int_{\Omega}\mathcal{H}(x,|u|)dx=\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx.$
The Musielak-Orlicz space $L^{\mathcal{H}}(\Omega)$, is defined by
$L^{\mathcal{H}}(\Omega)=\left\{u:\Omega\to\mathbb{R}\,\,\text{measurable};\,\,\rho_{\mathcal{H}}(u)<+\infty\right\},$
endowed with the Luxemburg norm
$\|u\|_{\mathcal{H}}=\inf\left\{\hat{\lambda}>0:\rho_{\mathcal{H}}\left(\frac{u}{\hat{\lambda}}\right)\leq 1\right\}.$
Analogous to Proposition 2.2, there are similar relationship between the modular $\rho_{\mathcal{H}}(\cdot)$ and the norm $\|\cdot\|_{\mathcal{H}}$, see [17, Proposition 2.13] for a detailed proof.
Proposition 2.6.
Assume $(H_{1})$ hold, $u\in L^{\mathcal{H}}(\Omega)$ and $\hat{\lambda}\in\mathbb{R}$. Then
  • $(i)$
    If $u\neq 0$, then $\|u\|_{\mathcal{H}}=\hat{\lambda}\Leftrightarrow\rho_{\mathcal{H}}(\frac{u}{\hat{\lambda}})=1$,
  • $(ii)$
    $\|u\|_{\mathcal{H}}<1\ (\text{resp.}\ >1,=1)\Leftrightarrow\rho_{\mathcal{H}}(\frac{u}{\hat{\lambda}})<1\ (\text{resp.}\ >1,=1)$,
  • $(iii)$
    If $\|u\|_{\mathcal{H}}<1\Rightarrow\|u\|_{\mathcal{H}}^{q^{+}}\leq\rho_{\mathcal{H}}(u)\leq\|u\|_{\mathcal{H}}^{p^{-}}$,
  • $(iv)$
    If $\|u\|_{\mathcal{H}}>1\Rightarrow\|u\|_{\mathcal{H}}^{p^{-}}\leq\rho_{\mathcal{H}}(u)\leq\|u\|_{\mathcal{H}}^{q^{+}}$,
  • $(v)$
    $\|u\|_{\mathcal{H}}\to 0\Leftrightarrow\rho_{\mathcal{H}}(u)\to 0$,
  • $(vi)$
    $\|u\|_{\mathcal{H}}\to+\infty\Leftrightarrow\rho_{\mathcal{H}}(u)\to+\infty$,
  • $(vii)$
    $\|u\|_{\mathcal{H}}\to 1\Leftrightarrow\rho_{\mathcal{H}}(u)\to 1$,
  • $(viii)$
    If $u_{n}\to u$ in $L^{\mathcal{H}}(\Omega)$, then $\rho_{\mathcal{H}}(u_{n})\to\rho_{\mathcal{H}}(u)$.
The Musielak-Orlicz Sobolev space $W^{1,\mathcal{H}}(\Omega)$ is defined by
$W^{1,\mathcal{H}}(\Omega)=\left\{u\in L^{\mathcal{H}}(\Omega):|\nabla u|\in L^{\mathcal{H}}(\Omega)\right\},$
and equipped with the norm
$\|u\|_{1,\mathcal{H}}=\|\nabla u\|_{\mathcal{H}}+\|u\|_{\mathcal{H}},$
where $\|\nabla u\|_{\mathcal{H}}=\|\,|\nabla u|\,\|_{\mathcal{H}}$.
The space $W_{0}^{1,\mathcal{H}}(\Omega)$ is defined as the closure of $C_{0}^{\infty}(\Omega)$ in $W^{1,\mathcal{H})}(\Omega)$. Note also that $L^{\mathcal{H}}(\Omega),W^{1,\mathcal{H}}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)$ are reflexive Banach spaces [17, Proposition 2.12].
We now present the following embedding relations given in [17, Proposition 2.16].
Proposition 2.7.
Assume that $(H_{1})$ and $(H_{2})$ hold. Then the following embeddings hold:
  • $(i)$
    $L^{\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega),W^{1,\mathcal{H}}(\Omega)\hookrightarrow W^{1,h(x)}(\Omega)$, $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow W_{0}^{1,h(x)}(\Omega)$ are continuous for $h\in C(\overline{\Omega})$ with $1\leq h(x)\leq p(x)$ for all $x\in\overline{\Omega}$.
  • $(ii)$
    $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ and $W_{0}^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{h(x)}(\Omega)$ are compact for $h\in C(\overline{\Omega})$ with $1\leq h(x)<p^{*}(x)$ for all $x\in\overline{\Omega}$.
Proposition 2.8.
[17] Assume that $(H_{1})$ and $(H_{2})$ hold. Then the following hold:
  • $(i)$
    The embedding $W^{1,\mathcal{H}}(\Omega)\hookrightarrow L^{\mathcal{H}}(\Omega)$ is compact;
  • $(ii)$
    There exists a constant $c>0$ independent of $u$ such that
    $\|u\|_{\mathcal{H}}\leq c\|\nabla u\|_{\mathcal{H}},\quad\forall u\in W_{0}^{1,\mathcal{H}}(\Omega).$
As a conclusion of Proposition 2.8, the space $W_{0}^{1,\mathcal{H}}(\Omega)$ can be equipped with an equivalent norm $\|\cdot\|_{1,\mathcal{H},0}$ given by
$\|u\|_{1,\mathcal{H},0}=\|\nabla u\|_{\mathcal{H}},\quad\forall u\in W_{0}^{1,\mathcal{H}}(\Omega).$
Proposition 2.9.
The convex functional
$\varrho_{\mathcal{H}}(u):=\int_{\Omega}\left(\frac{|\nabla u|^{p(x)}}{p(x)}+\mu(x)\frac{|\nabla u|^{q(x)}}{q(x)}\right)dx$
is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$, and its derivative $\varrho^{\prime}_{\mathcal{H}}$ satisfies the ($S_{+}$)-property [35] with the derivative
$\langle\varrho^{\prime}_{\mathcal{H}}(u),\varphi\rangle=\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla\varphi dx,$
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, where $\langle\cdot,\cdot\rangle$ is the dual pairing between $W_{0}^{1,\mathcal{H}}(\Omega)$ and its dual $W_{0}^{1,\mathcal{H}}(\Omega)^{*}$ [17].
Next, we present and prove a fractional-Hardy-type inequality which we use to keep the singular integral $\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$ under the control so that it does not exhibit an uncontrolled growth (blow up). When $\alpha\equiv 1$, the multi-phase case of (1) is given in [3], while the anisotropic case is given in [4].
Lemma 2.10.
Assume that $(H_{1}),(H_{2})$ hold. If, additionally, $\alpha p^{-}>1$ and $\alpha q^{+}>1$, then the following hold:
$(i)$ There exist constants $\hat{H},\kappa$ independent of $u$ such that the inequality
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx\leq\hat{H}\|u\|_{1,\mathcal{H},0}^{\kappa}$ (1)
holds for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$.
$(ii)$ There exists a constant $M>1$ such that $|x|<M$ for all $x\in\Omega\setminus\{0\}$. Then,
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx>\frac{1}{M^{\alpha q^{+}}}\|u\|^{{p^{-}\wedge q^{+}}}_{\mathcal{H}},\,\,\ \forall u\in W_{0}^{1,\mathcal{H}}(\Omega),$ (2)
where $p^{-}\wedge q^{+}=\min\{p^{-},q^{+}\}$ determined according to the magnitude of $u$.
Proof.
$(i)$ First, we obtain a fractional-Hardy-type inequality. To do so, we consider a general constant exponent $s$ with $1<s<N$.
For $u\in C_{0}^{\infty}(\Omega)$, we have
$\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)=\frac{N-\alpha s}{|x|^{\alpha s}}.$ (3)
Hence,
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx=\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s}\,\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)dx.$ (4)
Applying the Divergence Theorem along with integrating by parts provides
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx$ $=\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s}\,\nabla\cdot\left(\frac{x}{|x|^{\alpha s}}\right)dx$
$=-\frac{1}{N-\alpha s}\int_{\Omega}\mu(x)\,\nabla(|u|^{s})\cdot\frac{x}{|x|^{\alpha s}}dx,$ (5)
which is equal to
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx=-\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)\,|u|^{s-2}u\nabla u\cdot\frac{x}{|x|^{\alpha s}}dx.$ (6)
Employing the Cauchy-Schwarz inequality gives
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx$ $\leq\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)\,\frac{|u|^{s-1}}{|x|^{\alpha s-1}}|\nabla u|dx$
$=\frac{s}{N-\alpha s}\int_{\Omega}\mu(x)^{\frac{s-1}{s}}\,\frac{|u|^{s-1}}{|x|^{\alpha s-1}}\mu(x)^{\frac{1}{s}}\,|\nabla u|dx$ (7)
Applying Hölder’s inequality, and noticing that $(\alpha s-1)(\frac{s}{s-1})<\alpha s$, and assuming that $|x|<1$ (otherwise the result is trivial), gives
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\leq$ $\frac{s}{N-\alpha s}\left(\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{(\alpha s-1)(\frac{s}{s-1})}}dx\right)^{\frac{s-1}{s}}\left(\int_{\Omega}\mu(x)\,|\nabla u|^{s}dx\right)^{\frac{1}{s}}$
$\leq\frac{s}{N-\alpha s}\left(\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\right)^{\frac{s-1}{s}}\left(\int_{\Omega}\mu(x)\,|\nabla u|^{s}dx\right)^{\frac{1}{s}},$ (8)
which yields the following fractional-Hardy-type inequality
$\int_{\Omega}\mu(x)\,\frac{|u|^{s}}{|x|^{\alpha s}}dx\leq H_{\mu,s}\int_{\Omega}|\nabla u|^{s}dx,$ (9)
where $H_{\mu,s}:=\|\mu\|_{\infty}\left(\frac{s}{N-\alpha s}\right)^{s}$ (we will let $H_{1,s}:=\left(\frac{s}{N-\alpha s}\right)^{s}$ if $\mu\equiv 1$). However, by the density argument, (9) holds for any $u\in W_{0}^{1,s}(\Omega)$.
Next, to ensure a rigorous analysis without overlapping conditions, we partition $\Omega$ into the following disjoint measurable sets:
$\Omega_{<}^{(a)}:=\Omega_{<}\cap\{|u(x)|\geq|x|^{\alpha}\},\quad\Omega_{<}^{(b)}:=\Omega_{<}\cap\{|u(x)|<|x|^{\alpha}\},\quad$
$\Omega_{\geq}^{(c)}:=\Omega_{\geq}\cap\{|u(x)|<|x|^{\alpha}\},\quad\Omega_{\geq}^{(d)}:=\Omega_{\geq}\cap\{|u(x)|\geq|x|^{\alpha}\},$ (10)
where
$\Omega_{<}:=\{x\in\Omega:|\nabla u(x)|<1\},\quad\Omega_{\geq}:=\{x\in\Omega:|\nabla u(x)|\geq 1\}.$ (11)
Since $\Omega$ is bounded and connected, we obviously have $\Omega=\Omega_{<}^{(a)}\cup\Omega_{<}^{(b)}\cup\Omega_{\geq}^{(c)}\cup\Omega_{\geq}^{(d)}$.
Step 1: Analysis on $\Omega_{<}$.
For $x\in\Omega_{<}$, using (9) we have
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{q^{+}}dx\geq\frac{1}{H_{\mu,q^{+}}}\int_{\Omega_{<}}\mu(x)\,\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx,$ (12)
and
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{p^{+}}dx\geq\frac{1}{H_{1,p^{+}}}\int_{\Omega_{<}}\frac{|u|^{p^{+}}}{|x|^{\alpha p^{+}}}dx.$ (13)
Estimate on $\Omega_{<}^{(a)}$: In this region we have $|u(x)|\geq|x|^{\alpha}$, and hence
$\int_{\Omega_{<}^{(a)}}\mu(x)\,\frac{|u|^{q^{+}}}{|x|^{\alpha q^{+}}}dx\geq\int_{\Omega_{<}^{(a)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (14)
and
$\int_{\Omega_{<}^{(a)}}\frac{|u|^{p^{+}}}{|x|^{\alpha p^{+}}}dx\geq\int_{\Omega_{<}^{(a)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (15)
Therefore,
$\int_{\Omega_{<}^{(a)}}|\nabla u|dx\geq\frac{1}{2H_{1}}\int_{\Omega_{<}^{(a)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (16)
where $\frac{1}{H_{1}}:=\frac{1}{H_{1,p^{+}}}\wedge\frac{1}{H_{\mu,q^{+}}}$.
Estimate on $\Omega_{<}^{(b)}$: In this region we have $|u(x)|<|x|^{\alpha}$, therefore
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{q^{-}}dx\geq\frac{1}{H_{\mu,q^{-}}}\int_{\Omega_{<}}\mu(x)\,\frac{|u|^{q^{-}}}{|x|^{\alpha q^{-}}}dx,$ (17)
and
$\int_{\Omega_{<}}|\nabla u|dx\geq\int_{\Omega_{<}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{<}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx.$ (18)
Since
$\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q^{-}}}{|x|^{\alpha q^{-}}}dx\geq\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (19)
and
$\int_{\Omega_{<}^{(b)}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{<}^{(b)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (20)
we have
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{H_{\mu,q^{-}}}\int_{\Omega_{<}^{(b)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (21)
and
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{<}^{(b)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx.$ (22)
Therefore,
$\int_{\Omega_{<}^{(b)}}|\nabla u|dx\geq\frac{1}{2H_{2}}\int_{\Omega_{<}^{(b)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (23)
where $\frac{1}{H_{2}}:=\frac{1}{H_{1,p^{-}}}\wedge\frac{1}{H_{\mu,q^{-}}}$.
Step 2: Analysis on $\Omega_{\geq}$. For $x\in\Omega_{\geq}$, using (9) we have
$\int_{\Omega_{\geq}}|\nabla u|^{p(x)}dx\geq\int_{\Omega_{\geq}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{\mu,p^{-}}}\int_{\Omega_{\geq}}\mu(x)\,\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx,$ (24)
and
$\int_{\Omega_{\geq}}|\nabla u|^{p(x)}dx\geq\int_{\Omega_{\geq}}|\nabla u|^{p^{-}}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{\geq}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx,$ (25)
Estimate on $\Omega_{\geq}^{(c)}$: In this region we have $|u(x)|<|x|^{\alpha}$. Since
$\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (26)
and
$\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p^{-}}}{|x|^{\alpha p^{-}}}dx\geq\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx,$ (27)
it reads
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{H_{\mu,p^{-}}}\int_{\Omega_{\geq}^{(c)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx,$ (28)
and
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{H_{1,p^{-}}}\int_{\Omega_{\geq}^{(c)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx.$ (29)
Therefore,
$\int_{\Omega_{\geq}^{(c)}}|\nabla u|^{p(x)}dx\geq\frac{1}{2H_{3}}\int_{\Omega_{\geq}^{(c)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx,$ (30)
where $\frac{1}{H_{3}}:=\frac{1}{H_{1,p^{-}}}\wedge\frac{1}{H_{\mu,p^{-}}}$.
Estimate on $\Omega_{\geq}^{(d)}$: In this region, we have $|u(x)|\geq|x|^{\alpha}$. Considering the sufficiently large values of $|x|$, we have
$\int_{\Omega_{\geq}^{(d)}}\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}dx\leq\|\mu\|_{\infty}\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx,$ (31)
and
$\int_{\Omega_{\geq}^{(d)}}\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}dx\leq\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx.$ (32)
Hence,
$\int_{\Omega_{\geq}^{(d)}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\leq(1+\|\mu\|_{\infty})\int_{\Omega_{\geq}^{(d)}}|u|^{q^{+}}dx.$ (33)
Since $\Omega_{<}^{(a)},\Omega_{<}^{(b)},\Omega_{\geq}^{(c)}$ and $\Omega_{\geq}^{(d)}$ are measurable subsets of $\Omega$ with $\Omega=\Omega_{<}^{(a)}\cup\Omega_{<}^{(b)}\cup\Omega_{\geq}^{(c)}\cup\Omega_{\geq}^{(d)}$, we may extend all integrands by zero outside their respective sets in (16), (23), (30), and then integrate over all of $\Omega$, which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$
$\leq 2(H_{1}+H_{2})\int_{\Omega}|\nabla u|dx+2H_{3}\int_{\Omega}|\nabla u|^{p(x)}dx+(1+\|\mu\|_{\infty})\int_{\Omega}|u|^{q^{+}}dx.$ (34)
To conclude part $(i)$, we employ the embeddings in Proposition 2.7, which provides the desired result
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\,\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\leq\hat{H}\|u\|_{1,\mathcal{H},0}^{\kappa},$
where
$\kappa=\begin{cases}1&\text{if }\,\|u\|_{1,\mathcal{H},0}<1,\\ q^{+}&\text{if }\,\|u\|_{1,\mathcal{H},0}\geq 1,\end{cases}$
and $\hat{H}:=2(H_{1}+H_{2})\hat{c}_{1}\vee 2H_{3}\hat{c}_{2}\vee(1+\|\mu\|_{\infty})\hat{c}_{3}$, where $\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}$ are the corresponding embedding constants.
$(ii)$ Since $\Omega$ is bounded, there exists $M>1$ such that $|x|<M$ for all $x\in\Omega$. Thus, for any $x\in\Omega\setminus\{0\}$, we have $|x|^{\alpha p(x)}<M^{\alpha p(x)}\leq M^{\alpha p^{+}}$ and $|x|^{\alpha q(x)}<M^{\alpha q(x)}\leq M^{\alpha q^{+}}$. Therefore, $|x|^{\alpha p(x)}\vee|x|^{\alpha q(x)}<M^{\alpha q^{+}}$, which gives
$\int_{\Omega}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$ $>\frac{1}{M^{\alpha q^{+}}}\int_{\Omega}\left(|u|^{p(x)}+\mu(x)|u|^{q(x)}\right)dx$
$>\frac{1}{M^{\alpha q^{+}}}\|u\|^{{p^{-}\wedge q^{+}}}_{\mathcal{H}}.$ (35)
3 Variational Framework and Main Results
We define the energy functional $\mathcal{I}:W_{0}^{1,\mathcal{H}}(\Omega)\rightarrow\mathbb{R}$ corresponding to equation ($\mathcal{P}_{\lambda}$) by
$\mathcal{I}(u)=$ $\int_{\Omega}\left(\frac{|\nabla u|^{p(x)}}{p(x)}+\mu(x)\frac{|\nabla u|^{q(x)}}{q(x)}\right)dx+\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$-\lambda\int_{\Omega}F(x,u)dx,$
or
$\mathcal{I}(u)=\varrho_{\mathcal{H}}(u)+\mathcal{G}(u)-\lambda\int_{\Omega}F(x,u)dx,$
where $\mathcal{G}(u):=\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$.
Definition 3.1.
A function $u\in W_{0}^{1,\mathcal{H}}(\Omega)$ is called a (weak) solution to problem ($\mathcal{P}_{\lambda}$) if
$\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla\varphi dx+\int_{\Omega}\left(\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}\right)\varphi dx$
$=\lambda\int_{\Omega}f(x,u)\varphi dx,\,\,\ \forall\varphi\in W_{0}^{1,\mathcal{H}}(\Omega),$ (36)
where $F(x,t)=\int_{0}^{t}f(x,s)ds$.
Lemma 3.2.
$\mathcal{I}\in C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$ with the derivative
$\langle\mathcal{I}^{\prime}(u),\varphi\rangle$ $=\langle\varrho^{\prime}_{\mathcal{H}}(u),\varphi\rangle+\langle\mathcal{G}^{\prime}(u),\varphi\rangle-\lambda\langle f(x,u),\varphi\rangle,\,\,\ \forall\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$ (37)
Moreover, the critical points of $\mathcal{I}$ are the solutions of problem ($\mathcal{P}_{\lambda}$).
Proof.
By Proposition 2.9, $\varrho_{\mathcal{H}}$ is a continuously Gâteaux differentiable functional with the derivative $\langle\varrho^{\prime}_{\mathcal{H}}(\cdot),\varphi\rangle$. Moreover, using the assumption $(\mathbf{f}_{1})$, and applying Holder inequality along with the related embeddings shows that the functional $\int_{\Omega}F(x,u)dx$ is also continuously Gâteaux differentiable with the derivative $\langle f(x,u),\varphi\rangle=\int_{\Omega}f(x,u)\varphi dx$. Thus, it remains only to determine the Gâteaux derivative of $\mathcal{G}$ and show that it has the same regularity on $W_{0}^{1,\mathcal{H}}(\Omega)$.
To begin, by the Mean Value Theorem, there are $0<\varepsilon_{1},\varepsilon_{2}<1$ such that
$\langle\mathcal{G}^{\prime}(u),\varphi\rangle$ $=\lim_{t\to 0}\int_{\Omega}\frac{1}{t}\left(\frac{|u+t\varphi|^{p(x)}-|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{\mu(x)|u+t\varphi|^{q(x)}-|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$=\lim_{t\to 0}\int_{\Omega}\frac{1}{t}\left(\frac{d}{d\gamma}\frac{|u+\gamma t\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\bigg|_{\gamma=\varepsilon_{1}}+\frac{d}{d\gamma}\frac{\mu(x)|u+\gamma t\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\bigg|_{\gamma=\varepsilon_{2}}\right)dx$
$=\lim_{t\to 0}\int_{\Omega}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\right$.
$+\left.|x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\right)dx,$
for all $u,\varphi\in W_{0}^{1,\mathcal{H}}(\Omega)$, $\gamma\in\mathbb{R}$. Using the Young inequality, we obtain
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|$ $\leq k_{1}|x|^{-\alpha p(x)}(|u|^{p(x)}+|\varphi|^{p(x)}),$ (39)
and
$\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$ $\leq k_{2}|x|^{-\alpha q(x)}(\mu(x)|u|^{q(x)}+\mu(x)|\varphi|^{q(x)}).$ (40)
where $k_{1}:=\frac{2^{p(x)-1}(p(x)-1)+1}{p(x)}$, and $k_{2}:=\frac{2^{q(x)-1}(q(x)-1)+1}{q(x)}$.
Using (39),(40) and Lemma 2.10, we can write
$\bigg||x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\varphi\bigg|+\bigg||x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\varphi\bigg|$
$\leq k_{1}(|x|^{-\alpha p(x)}|u|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|u|^{q(x)})+k_{2}(|x|^{-\alpha p(x)}|\varphi|^{p(x)}+|x|^{-\alpha q(x)}\mu(x)|\varphi|^{q(x)})$
$\leq k\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right),$ (41)
where $k:=\frac{2^{q^{+}}(q^{+}-1)}{p^{-}}\vee k_{1}\vee k_{2}$.
Therefore, by the Lebesgue Dominated Convergence Theorem it reads
$\langle\mathcal{G}^{\prime}(u),\varphi\rangle$ $=\int_{\Omega}\lim_{t\to 0}\left(|x|^{-\alpha p(x)}|u+\varepsilon_{1}t\varphi|^{p(x)-2}(u+\varepsilon_{1}t\varphi)\right$.
$+\left.|x|^{-\alpha q(x)}\mu(x)|u+\varepsilon_{2}t\varphi|^{q(x)-2}(u+\varepsilon_{2}t\varphi)\right)\varphi dx,$
$=\int_{\Omega}\left(|x|^{-\alpha p(x)}|u|^{p(x)-2}u+\mu(x)|x|^{-\alpha q(x)}|u|^{q(x)-2}u\right)\varphi dx.$ (42)
Since the right-hand side of (3), as a function of $\varphi$, is a linear functional, $\mathcal{G}^{\prime}$ is linear on $W_{0}^{1,\mathcal{H}}(\Omega)$.
Next, by the Young inequality and Lemma 2.10, we obtain
$|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|$ $\leq\int_{\Omega}\left(|x|^{-p(x)}|u|^{p(x)-1}|\varphi|+|x|^{-\alpha q(x)}\mu(x)|u|^{q(x)-1}|\varphi|\right)dx$
$\leq\frac{q^{+}}{p^{-}}\int_{\Omega}\left(\frac{|u|^{p(x)}}{p(x)|x|^{\alpha p(x)}}+\frac{|\varphi|^{p(x)}}{p(x)|x|^{\alpha p(x)}}\right)+\left(\frac{\mu(x)|u|^{q(x)}}{q(x)|x|^{\alpha q(x)}}+\frac{\mu(x)|\varphi|^{q(x)}}{q(x)|x|^{\alpha q(x)}}\right)dx$
$\leq\frac{q^{+}}{p^{-}}\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+\|\varphi\|_{1,\mathcal{H},0}^{\kappa}\right).$ (43)
Therefore, for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$, we have
$\|\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}$ $=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}|\langle\mathcal{G}^{\prime}(u),\varphi\rangle|\leq\frac{q^{+}}{p^{-}}\hat{H}\left(\|u\|_{1,\mathcal{H},0}^{\kappa}+1\right),$ (44)
which means that $\mathcal{G}^{\prime}$ is bounded. Therefore, $\mathcal{G}$ is Gâteaux differentiable whose derivative is given by the formula (3). Now, we proceed with the continuity of $\mathcal{G}^{\prime}$. Let $(u_{n})\subset W_{0}^{1,\mathcal{H}}(\Omega)$ such that $u_{n}\to u$ in $W_{0}^{1,\mathcal{H}}(\Omega)$. Applying the inequality (Proposition 17.2, [14])
$\left|\left|\xi\right|^{m-2}\xi-\left|\psi\right|^{m-2}\psi\right|\leq C_{m}\left|\xi-\psi\right|\{|\xi|+|\psi|\}^{m-2},\quad\xi,\psi\in\mathbb{R}^{N},\,\,m>1,$ (45)
gives
$\left|\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u),\varphi\rangle\right|$ $\leq\int_{\Omega}|x|^{-p(x)}\left||u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right||\varphi|dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left||u_{n}|^{q(x)-2}u_{n}-|u|^{q(x)-2}u\right||\varphi|dx$
$\leq C_{p}\int_{\Omega}|x|^{-p(x)}\{|u_{n}|+|u|\}^{p(x)-2}|u_{n}-u||\varphi|dx$
$+C_{q}\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\{|u_{n}|+|u|\}^{q(x)-2}|u_{n}-u||\varphi|dx.$ (46)
Thus, considering that $u_{n}\to u$ in $L(\Omega)$ and $(u_{n})$ is bounded, it reads
$\|\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u)\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}=\sup_{\|\varphi\|_{1,\mathcal{H},0}\leq 1}\left|\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u),\varphi\rangle\right|\to 0.$ (47)
Therefore, $\mathcal{G}$ is of class $C^{1}(W_{0}^{1,\mathcal{H}}(\Omega),\mathbb{R})$.
Finally, $\mathcal{I}^{\prime}(u)=0$ is exactly the weak formulation of ($\mathcal{P}_{\lambda}$), so critical points of $\mathcal{I}$ are weak solutions. This completes the proof. ∎
The main results of the present paper is given below.
Theorem 3.3.
Assume that the following assumptions hold:
  • $(\mathbf{\beta_{0}})$
    $\beta\in C_{+}\left(\overline{\Omega}\right)$ such that $q^{+}<\beta^{-}\leq\beta^{+}<p^{\ast}(x)$   $\forall x\in\overline{\Omega}$.
  • $(\mathbf{f}_{1})$
    $f:\overline{\Omega}\times\mathbb{R}\rightarrow\mathbb{R}$ is a Carathéodory function and satisfies the growth condition
    $\left|f(x,t)\right|\leq c_{1}+c_{2}\left|t\right|^{\beta(x)-1},\quad\forall\left(x,t\right)\in\overline{\Omega}\times\mathbb{R},$
    where $c_{1}$ and $c_{2}$ are positive constants.
  • $(\mathbf{f}_{2})$
    $f(x,t)=o\left(\left|t\right|^{q^{+}-1}\right)$, $t\rightarrow 0$ uniformly $\forall x\in\overline{\Omega}$.
  • $(\mathbf{AR})$
    $\exists K>0$, $\theta>q^{+}$ such that
    $0<\theta F(x,t)\leq f(x,t)t,\quad\left|t\right|\geq K\text{ a.e.}\,x\in\overline{\Omega}.$
Then problem ($\mathcal{P}_{\lambda}$) has at least one nontrivial positive weak solution.
Note that $(\mathbf{AR})$ stands for the Ambrosetti–Rabinowitz condition [2].
To obtain the result of Theorem 3.3, we need to show Lemma 3.4 and Lemma 3.6 hold.
Lemma 3.4.
Suppose $(\mathbf{\beta_{0}})$, $(\mathbf{f}_{1})$, $(\mathbf{f}_{2})$ and $(\mathbf{AR})$ hold. Then the following statements hold:
  • $(i)$
    There exist two positive real numbers $\gamma$ and $\eta$ such that $\mathcal{I}(u)\geq\eta>0$, for all $u\in W_{0}^{1,\mathcal{H}}(\Omega)$ with $\|u\|_{1,\mathcal{H},0}=\gamma$.
  • $(ii)$
    There exists a $\hat{u}\in W_{0}^{1,\mathcal{H}}(\Omega)$ such that $\|\hat{u}\|_{1,\mathcal{H},0}>\gamma$  and   $\mathcal{I}(\hat{u})<0$.
Proof.
$(i)$ By $(\mathbf{f}_{1})$ and $(\mathbf{f}_{2})$, one can write
$F(x,t)\leq\varepsilon|t|^{q^{+}}+C_{\varepsilon}|t|^{\beta(x)},\quad\forall\left(x,t\right)\in\overline{\Omega}\times\mathbb{R}.$ (48)
Then
$\mathcal{I}(u)$ $\geq\frac{1}{q^{+}}\rho_{\mathcal{H}}(\nabla u)+\frac{1}{q^{+}M^{\alpha\tau}}\rho_{\mathcal{H}}(u)-\lambda\varepsilon\int_{\Omega}|u|^{q^{+}}dx-\lambda C_{\varepsilon}\int_{\Omega}|u|^{\beta(x)}dx$
$\geq\frac{1}{q^{+}}\rho_{\mathcal{H}}(\nabla u)-\lambda\varepsilon|u|^{q^{+}}_{q^{+}}-\lambda C_{\varepsilon}\left(|u|^{\beta^{+}}_{\beta^{+}}+|u|^{\beta^{-}}_{\beta^{-}}\right).$ (49)
Using Proposition 2.7, and assuming that $\|u\|_{1,\mathcal{H},0}<1$ provides
$\mathcal{I}(u)$ $\geq\frac{1}{q^{+}}\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda\varepsilon c_{3}\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda C_{\varepsilon}c_{4}\|u\|_{1,\mathcal{H},0}^{\beta^{-}}$
$=\left(\frac{1}{q^{+}}-\lambda\varepsilon c_{3}\right)\|u\|_{1,\mathcal{H},0}^{q^{+}}-\lambda C_{\varepsilon}c_{4}\|u\|_{1,\mathcal{H},0}^{\beta^{-}}.$ (50)
[…]
$(ii)$ Due to $(\mathbf{AR})$, we can write
$F(x,t)\geq c_{5}|t|^{\theta},\quad\forall x\in\overline{\Omega},\,\,|t|\geq K.$ (51)
Thus, for $0\neq\phi\in W_{0}^{1,\mathcal{H}}(\Omega)$ and $t>1$, we have
$\mathcal{I}(t\phi)$ $\leq\frac{t^{q^{+}}}{p^{-}}\left(\rho_{\mathcal{H}}(\nabla\phi)+\int_{\Omega}\left(\frac{|\phi|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|\phi|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx\right)-\lambda c_{5}t^{\theta}\int_{\Omega}|u|^{\theta}dx,$ (52)
which implies that $\mathcal{I}(t\phi)\to$-$\infty$ as $t\to\infty$. Therefore, letting $\hat{u}=t\phi$, and hence $\|\hat{u}\|_{1,\mathcal{H},0}=t\|\phi\|_{1,\mathcal{H},0}$, leads to $\|\hat{u}\|_{1,\mathcal{H},0}>1>\gamma$ and $\mathcal{I}(\hat{u})<0$ provided $t$ is large enough. ∎
Definition 3.5.
Let $X$ be a Banach space, and $J:X\rightarrow\mathbb{R}$ be a $C^{1}$-functional. We say that $J$ satisfies the Palais-Smale condition ($(\mathbf{PS})$ for short) [2] if: every sequence $(u_{n})\subset X$ such that $J(u_{n})$ is bounded and $J^{\prime}(u_{n})\rightarrow 0$ in $X^{*}$ admits a convergent subsequence in $X$.
Lemma 3.6.
Suppose $(\mathbf{\beta_{0}})$, $(\mathbf{f}_{1})$ and $(\mathbf{AR})$ hold. Then $\mathcal{I}$ satisfies the $(\mathbf{PS})$ condition.
Proof.
Let assume that there exists a sequence $(u_{n})\subset W_{0}^{1,\mathcal{H}}(\Omega)$ such that
$|\mathcal{I}(u_{n})|\leq c_{0}\quad\text{ and }\quad\|\mathcal{I}^{\prime}(u_{n})\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}\rightarrow 0,\quad 0<c_{0}\in\mathbb{R}.$ (53)
Using (53), $(\mathbf{AR})$, and Proposition 2.6, we obtain
$c_{0}$ $\geq\mathcal{I}(u_{n})=\varrho_{\mathcal{H}}(u_{n})+\mathcal{G}(u_{n})-\lambda\int_{\Omega}F(x,u_{n})dx$
$\geq\left(\frac{1}{q^{+}}-\frac{1}{\theta}\right)\rho_{\mathcal{H}}(\nabla u_{n})+\frac{1}{\theta}\int_{\Omega}\left(\rho_{\mathcal{H}}(\nabla u_{n})+\frac{|u_{n}|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u_{n}|^{q(x)}}{|x|^{\alpha q(x)}}-\lambda f(x,u_{n})u_{n}\right)dx$
$\geq\left(\frac{1}{q^{+}}-\frac{1}{\theta}\right)\|u_{n}\|_{1,\mathcal{H},0}^{p^{-}}-\frac{1}{\theta}\|u_{n}\|_{1,\mathcal{H},0}\|\mathcal{I}^{\prime}(u_{n})\|_{W_{0}^{1,\mathcal{H}}(\Omega)^{*}}.$ (54)
Therefore $(u_{n})$ is bounded in $W_{0}^{1,\mathcal{H}}(\Omega)$. Since $W_{0}^{1,\mathcal{H}}(\Omega)$ is reflexive, passing to a subsequence, we have $u_{n}\rightharpoonup u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$. Then using $(\mathbf{f}_{1})$, Hölder inequality, and Proposition 2.7, we obtain
$\left|\langle f(x,u_{n}),u_{n}-u_{0}\rangle\right|\leq c_{1}\left||u_{n}|^{\beta(x)-1}\right|_{\frac{\beta(x)}{\beta(x)-1}}\left|u_{n}-u_{0}\right|_{\beta(x)}+c_{2}\int_{\Omega}\left|u_{n}-u_{0}\right|dx\to 0.$ (55)
First let’s write
$\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle=\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle+\langle\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle.$ (56)
Taking into consideration that $\mathcal{G}^{\prime}(u_{0})\in W^{1,\mathcal{H}}(\Omega)^{*}$, we obtain
$\langle\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle=\int_{\Omega}\mathcal{G}^{\prime}(u_{0})(u_{n}-u_{0})dx\to 0.$ (57)
Moreover, since $u_{n}\to u_{0}$ in $L^{2}(\Omega)$, $(u_{n})$ is bounded. Thus
$\langle\mathcal{G}^{\prime}(u_{n})-\mathcal{G}^{\prime}(u_{0}),u_{n}-u_{0}\rangle$ $=\int_{\Omega}|x|^{-\alpha p(x)}\left(|u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right)(u_{n}-u_{0})dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left(|u_{n}|^{q(x)-2}u_{n}-|u_{0}|^{q(x)-2}u_{0}\right)(u_{n}-u_{0})dx$
$\leq\int_{\Omega}|x|^{-\alpha p(x)}\left||u_{n}|^{p(x)-2}u_{n}-|u_{0}|^{p(x)-2}u_{0}\right||u_{n}-u_{0}|dx$
$+\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\left||u_{n}|^{q(x)-2}u_{n}-|u_{0}|^{q(x)-2}u_{0}\right||u_{n}-u_{0}|dx$
$\leq C_{p}\int_{\Omega}|x|^{-\alpha p(x)}\{|u_{n}|+|u_{0}|\}^{p(x)-2}|u_{n}-u_{0}|^{2}dx$
$+C_{q}\int_{\Omega}|x|^{-\alpha q(x)}\mu(x)\{|u_{n}|+|u_{0}|\}^{q(x)-2}|u_{n}-u_{0}|^{2}dx\to 0$ (58)
where we applied the inequality (45). Thus
$\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle\to 0.$ (59)
Now, using (53) gives
$\langle\mathcal{I}^{\prime}(u_{n}),u_{n}-u_{0}\rangle$ $=\langle\varrho^{\prime}_{\mathcal{H}}(u_{n}),u_{n}-u_{0}\rangle+\langle\mathcal{G}^{\prime}(u_{n}),u_{n}-u_{0}\rangle-\lambda\langle f(x,u_{n}),u_{n}-u_{0}\rangle\to 0.$ (60)
However, (55), (59) and (60) together implies that
$\langle\varrho^{\prime}_{\mathcal{H}}(u_{n}),u_{n}-u_{0}\rangle\to 0.$ (61)
Thus, by Proposition 2.9 $($$S_{+}$)-property), $u_{n}\to u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$. In conclusion, $\mathcal{I}$ satisfies the $(\mathbf{PS})$ condition. ∎
Proof of Theorem 3.3.
By Lemma 3.4, $\mathcal{I}$ has Mountain-Pass geometry; and by Lemma 3.6, $\mathcal{I}$ satisfies $(\mathbf{PS})$. Also, $\mathcal{I}(0)=0$. Thus, by the Mountain-Pass Theorem [34]:
  • $u_{0}$ is a critical point of $\mathcal{I}$ at mountain pass level $c_{MP}$:
    $c_{MP}:=\inf_{\gamma\in\Gamma}\max_{t\in[0,1]}\mathcal{I}(\gamma(t))\geq\eta>0,$
    where $\Gamma=\{\gamma\in C([0,1],W^{1,\mathcal{H}}_{0}(\Omega)):\gamma(0)=0,\gamma(1)=\hat{u}\}$.
  • Since $\mathcal{I}(u_{0})\geq\eta>0$, $u_{0}\neq 0$ is a nontrivial critical point of $\mathcal{I}$ at level $c_{MP}$;
    $\mathcal{I}(u_{0})=c_{MP}$ and $\mathcal{I}^{\prime}(u_{0})=0$.
Therefore, by Lemma 3.2, $u_{0}\in W_{0}^{1,\mathcal{H}}(\Omega)$ corresponds to a nontrivial weak solution of ($\mathcal{P}_{\lambda}$).
Define $f_{+}:\overline{\Omega}\times\mathbb{R}\rightarrow[0,\infty)$ by
$f_{+}(x,t)=\left\{\begin{array}[]{cc}f(x,t)&\text{if }\,t\geq 0,\\ 0&\text{if }\,t<0.\end{array}\right$.
Then clearly the modified energy functional $\mathcal{I_{+}}$ is also of class $C^{1}$ on $W_{0}^{1,\mathcal{H}}(\Omega)$.
Step 1.
Define $u^{-}:=\max\{-u,0\}$. We test (3.1) by replacing $u^{-}$ with $\varphi$ which gives
$\int_{\Omega}(|\nabla u|^{p(x)-2}\nabla u+\mu(x)|\nabla u|^{q(x)-2}\nabla u)\cdot\nabla u^{-}dx+\int_{\Omega}\left(\frac{|u|^{p(x)-2}u}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)-2}u}{|x|^{\alpha q(x)}}\right)u^{-}dx$
$=\lambda\int_{\Omega}f(x,u)u^{-}dx$ (62)
for a nontrivial weak solution $u$ of ($\mathcal{P}_{\lambda}$). Then considering the decomposition $\Omega=\{x\in\Omega:u(x)\geq 0\}\cup\{x\in\Omega:u(x)<0\}$ for (3), we have
$0\geq-\int_{\{u<0\}}(|\nabla u|^{p(x)}+\mu(x)|\nabla u|^{q(x)})dx-\int_{\{u<0\}}\left(\frac{|u|^{p(x)}}{|x|^{\alpha p(x)}}+\mu(x)\frac{|u|^{q(x)}}{|x|^{\alpha q(x)}}\right)dx$
$=\lambda\int_{\{u\geq 0\}}f(x,u)u^{-}dx\geq 0,$ (63)
which implies that $u^{-}\equiv 0$. So, since $u=0$ on $\partial\Omega$, by the weak minimum principle there is a set of positive measure where $u\geq 0$, and $u\not\equiv 0$.
Step 2.
We claim that $u(x)>0$ for all $x\in\Omega$. Suppose, on the contrary, that there exists an interior point $x_{0}\in\Omega$ such that $u(x_{0})=0$. Since $u\geq 0$ a.e., this must be a global minimum of $u$. This, by the strong minimum principle, implies that $u$ is constant. That is, $u\equiv 0$ in a neighborhood of $x_{0}$. However, this contradicts $u\not\equiv 0$. Therefore $u(x)>0$ for all $x\in\Omega$.
Funding
This work was supported by Athabasca University Research Incentive Account [140111 RIA].
ORCID
https://orcid.org/0000-0002-6001-627X